# Relative pose of three calibrated and partially calibrated cameras from four points using virtual correspondences

### v2

@@ Abstract
We study challenging problems of estimating the relative pose of three cameras and propose novel efficient solutions to (1) the notoriously difficult configuration of four points in three calibrated views, known as the 4p3v problem, and (2) to the previously unsolved configuration of four points in three cameras with unknown shared focal length, i.e., the 4p3vf problem. Our solutions are based on the simple idea of generating one or two additional virtual point correspondences in two views by using the information from the locations of the four input correspondences in the three views. We generate such correspondences using either a very simple and efficient strategy where the new points are the mean points of three corresponding input points or using a simple neural network. The new solvers are efficient and easy to implement since they are based on existing efficient minimal solvers, i.e., the well-known 5-point and 6-point relative pose solvers and the P3P solver. Our solvers achieve state-of-the-art results on real data.
@@ 1 Introduction
Camera geometry estimation is crucial in many computer vision applications, e.g., visual navigation [51], structure-from-motion [55], augmented and mixed reality [9], self-driving cars [18], and visual localization [50].
Due to noise and outliers in input correspondences, the predominant way in camera geometry estimation is to use a hypothesis-and-test framework, e.g., RANSAC [16, 11, 48, 3]. For RANSAC-like methods, it is critical to estimate camera geometry using as few correspondences as possible since the number of RANSAC iterations (and, thus, its run-time) depends exponentially on the number of correspondences required for the model estimation.
@@
Figure 1: Visualization of the 4p3v problem and our solution that is based on generating new virtual correspondence(s) (in red) between two views. This is done using coordinates of input point correspondences. Then the 4p3v problem is solved by existing efficient minimal solvers, i.e., the 5pt [38] and the P3P solvers [43].
Algorithms that solve camera geometry estimation problems using all available constraints and the minimum number of correspondences, such that the resulting system of equations has a finite number of solutions, are called minimal solvers, and the corresponding problems are called minimal problems. Minimal problems often result in complex systems of polynomial equations in several variables. After introducing algebraic-based methods [57, 26, 28, 30, 6] for generating efficient polynomial solvers to the computer vision community, solutions to many previously unsolved minimal problems were proposed
[58, 8, 36, 31, 27, 29, 60, 7]. Still, some problems result in equations for which state-of-the-art algebraic methods fail to generate a solver that is efficient and/or numerically stable.
@@
In this paper, we study such challenging problems of estimating the relative pose of three cameras. These problems have attracted attention for a long time [20, 47, 33, 37, 1]. However, due to their complexity, they are still not considered fully solved. No efficient and practical solutions exist for most of the minimal configurations of point and line correspondences, as well as for partially calibrated cameras. One such configuration that is particularly interesting is the notoriously difficult configuration of four points in three views [47, 38]. This configuration is minimal for three cameras with an unknown shared focal length, i.e., the 4p3vf problem, and provides one more constraint than minimal for calibrated cameras, i.e., the 4p3v problem.
@@
State-of-the-art algebraic and numerical methods are known to fail to generate efficient and numerically stable solutions to both the 4p3v and 4p3vf problems. The existing methods for solving the 4p3v problem for calibrated cameras are thus only approximate. By solving only for one (few) solutions from the 272 solutions of the system that describe the problem [21], and by discretely sampling the space of potential solutions [39], the existing 4p3v methods can often fail, i.e., the returned solution can be, in general, arbitrarily far from the geometrically correct solution. To decrease the failure rate, both methods [21, 39] require a lot of tuning and are not easy to re-implement. [^1] For the solver proposed in [39], there is no publicly available implementation. The publicly available implementation of the solver from [21] is quite complex and requires a non-negligible effort to run.
@@
In this paper, we propose a novel approach for solving both the calibrated 4p3v, as well as, the previously unsolved 4p3vf problem for cameras with an unknown shared focal length. Our solutions are based on the simple idea of generating new approximate point correspondence(s) between two of the three views. [^2] Note that similarly to the state-of-the-art solutions [39, 21], our solutions are only approximate. However, as we show in the experiments, they provide good initialization for LO-RANSAC and outperform [21]. Such correspondences are generated using only the locations of the four triplets of input point correspondences. Since the generation of new points does not require information from the image itself (e.g., information about appearance or features in the image), new correspondences, in general, do not need to correspond to any physical 3D points in the scene. Thus, we call them virtual correspondences. Using virtual correspondences, we can efficiently solve the 4p3v and 4p3vf problems by first estimating the relative pose of the two cameras from five/six correspondences using efficient 5pt/6pt solvers [38, 59], and then registering the third camera using the P3P solver [43]. We call this combination of the 5pt/6pt relative pose and the P3P solver the 5pt+P3P and 6pt+P3P solvers.
@@
Based on this idea, we propose two groups of novel solvers for the 4p3v and 4p3vf problems: (1) M-solvers (4p3v(M),4p3vf(M)): Solvers that use the mean points of three corresponding points detected in two views and potentially points in their vicinity ((M±$\pm\delta$)-solvers) as new virtual point correspondences. (2) L-solvers (4p3v(L),4p3vf(L)): Solvers that, given four triplets of corresponding points in three views and a point in the first view, use a network to predict a corresponding point to this point in the second view, i.e., a virtual correspondence. While simple to implement, the novel solvers achieve state-of-the-art results for the 4p3v/4p3vf problems on real data.
@@
  • [leftmargin=0.3cm]
  • For the well-known challenging 4p3v problem for calibrated cameras, we propose two groups of novel solutions 4p3v(M)- and 4p3v(L)-based solvers. These solvers generate an additional virtual point correspondence(s) in two views by leveraging the locations of four input triplet correspondences. The new solvers achieve state-of-the-art results in terms of accuracy on real data. Compared to state-of-the-art solutions to the 4p3v problem [21, 39], which are non-trivial and difficult to re-implement such that they are numerically stable and fast, our new solutions, especially the 4p3v(M)-based solvers, can be easily implemented using existing efficient implementations of the 5pt solver [38] and the P3P solver [43].
  • We provide the first efficient solutions to the previously unsolved 4p3vf problem for cameras with unknown equal focal lengths. Our novel 4p3vf(M)- and 4p3vf(L)-based solvers generate for each instance two virtual correspondences to solve the problem using the efficient 6pt [38] and the P3P solver [43]. Our novel solvers show the potential of the proposed idea of virtual correspondences to be applied to other camera geometry problems.
  • We compare the proposed solvers, the state-of-the-art 4p3v solver [21], as well as “oracle”-based 4p3v/4p3vf solvers, with the baseline minimal 5pt+P3P and 6pt+P3P solvers. By evaluating all solvers on a large amount of real data as well as data with outliers, we discuss which minimal point configuration is potentially the most useful in which scenario.
    The source code of our solvers will be publicly available.
@@ 2 Related work
For three uncalibrated cameras, 6 point correspondences are necessary to estimate the trifocal tensor, with a solution known for a long time [46, 61]. Solutions to three minimal combinations of points and lines are presented in [40]. The minimal configuration using 9 lines is much more challenging and was solved only recently by Larsson et al. [30]. However, the final solver is far from practical as it performs elimination of a huge $16k\times 13k$ matrix and runs 17.8s.
@@
For calibrated cameras, the configuration that attracts most of the attention is the configuration of four points in three views (the 4p3v problem). Note, that this is not a minimal configuration since it generates 12 constraints for 11 degrees-of-freedom (DoF) (see also Section 3). The 4p3v problem is known as being extremely difficult to solve. Several papers present mostly theoretical results [33, 37, 1]. For four triplets of exact points without noise, it is shown that the 4p3v problem has, in general, a unique solution [20, 47].
@@
To the best of our knowledge, there are only two reasonably efficient solutions to the 4p3v problem reported in the literature. The first solver [39] is based on one-dimensional exhaustive search. In the paper, the authors derive several interesting theoretical results and show that the four point correspondences between two calibrated views constrain the epipole in each image to lie on a curve of degree ten. The solver performs a one-dimensional sweep of the curve of possible epipoles. For each potential epipole, it computes the relative pose of two cameras, registers the third camera using three triangulated points, and finally extracts the solution minimizing the reprojection error of the fourth point in the third view. The evaluation of the solver on one potential epipole is fast. Yet, to obtain reasonable precise and stable results, usually, 1,000 candidates need to be evaluated and even then, refinement at multiple local minima is required to improve the precision. The reported run-times of this solver were $1-12$ms depending on the number of searched points. [^3] This is the runtime reported in the paper from 2006, however, implemented by a highly skilled researcher (the author of the fastest version of the well-known 5pt relative pose solver). Therefore, we do not expect significant speedup on recent hardware. There is no publicly available implementation for this solver and it is not easy to re-implement. As such, the literature does not compare against the solver in experiments. As an upper bound of the performance of [39], we compare against an oracle version using the true epipole.
@@
The second efficient solver to the 4p3v problem was published only recently [21]. In this paper, the authors first transform the 4p3v problem to a minimal problem by considering a line passing through the last correspondence in the third view. The resulting system of equations is solved using an efficient Homotopy continuation (HC) method [15, 56]. To avoid computing large numbers of spurious solutions, an MLP-based classifier is trained. For a given problem $p$, it selects one or several starting problem-solution pairs (so-called anchors), such that the geometrically meaningful/correct solution of $p$ can be obtained by HC starting from this anchor. This strategy is fast, running $16.3\mu s$ on average per solution. At the same time, it has a high failure rate. The success rate of the 4p3v solver reported in [21] on two test datasets and data without noise is $26.3\%$. [21] do not show results of their solver in a real scenario, i.e., a RANSAC-like framework with noisy data. We provide such an evaluation in this paper.
@@
Simple solutions to the 4p3v problem for orthographic and scaled orthographic views were presented in [63, 34]. In [34], the author suggested an iterative approach for updating to perspective views. However, he reported results only on a few synthetic instances. Even after spending a non-negligible amount of work and time (months) to make this update work, we were not able to get reasonable results on real data with general perspective cameras.
@@
Minimal configurations of points and lines in three calibrated views were studied in [14, 24, 13], aiming to classify and derive the number of solutions for different configurations. Solutions to two minimal configurations of combinations of points and lines, the so-called Chicago and Cleveland problems, were proposed in [15]. These problems result in 216 respectively 312 solutions and are solved using HC methods [56]. Due to their complexity, the resulting solvers are not practical in real applications, with running times ranging from $0.66$s to $1.9$s.
Recently, a very efficient GPU implementation of the homotopy continuation method was used to solve minimal problems of four points in three views and six lines in three views for a generalized (non-central) camera [12].
@@
In our solutions, we generate virtual correspondences. Virtual matches are also used in the literature of affine correspondences (AC) [42, 44, 45, 5]. There, points are sampled based on the affine feature geometry to generate point correspondences from affine ones. In our scenario, we are directly given point correspondences, without associated feature geometry, and we predict additional point matches.
@@ 3 Estimating the relative pose of three cameras
Let $N$ cameras observe a set of 3D points $\mathcal{P}$ and for each point $\mathbf{P}\in\mathcal{P}$, let $C_{\mathbf{P}}>1$ be the number of cameras that observe it. A necessary condition for a relative pose problem of $N$ calibrated cameras to be minimal is [15]
@@
A configuration of points in $N=3$ views that satisfies the constraint (1) of a minimal problem is three points visible in all three cameras and two additional points visible in two of the three cameras. We will call this configuration $(5_{3},5_{3},3_{3})$, where the lower index on the position $i$ indicates how many from the image points sampled in the camera $i$ are visible in all three cameras.
The configuration of four points visible in all three cameras, i.e., the configuration $(4_{4},4_{4},4_{4})$, generates an over-constrained problem. In this case, we have one more constraint than DoF, i.e., in Eq. (1), we have $12>11$. A minimal solution would need to drop one constraint, e.g., by considering only a line passing through one of the points in the third view [21] or by considering a “half” point correspondence. Since, in practice, we always have full correspondences and sampling one less point in one view leads to an under-constrained problem, the configuration $(4_{4},4_{4},4_{4})$, is usually considered “minimal”.
For cameras with unknown common focal lengths, we have one more DoF. This means that, for $N=3$, the right-hand side of equation (1) equals to 12, resulting in $(6_{3},6_{3},3_{3})$ and $(4_{4},4_{4},4_{4})$ being minimal configurations.
@@ 3.1 Calibrated cameras
In this section, we describe solutions for three calibrated cameras. We start with one baseline solution for the minimal $(5_{3},5_{3},3_{3})$ configuration followed by novel solutions for the “minimal” $(4_{4},4_{4},4_{4})$ configuration.
5pt+P3P solver:  First, we review a simple baseline minimal solver for the configuration $(5_{3},5_{3},3_{3})$, i.e., the configuration where three points are visible in all three cameras and two additional points in two of the three cameras.
@@
The 5pt+P3P solver first estimates the relative pose of two cameras from 5 image point correspondences using the efficient 5pt solver [38]. Next, the three points that are visible in all three views are triangulated. Finally, the third camera is registered using the three 2D-3D point correspondences and the well-known efficient P3P solver [43].
@@
This solver is straightforward and it is based on efficient existing solvers [38, 43]. The 5pt+P3P solver is not novel. This solver and the $(5_{3},5_{3},3_{3})$ configuration was, e.g., mentioned in [14], where the authors present a classification of minimal problems for points and lines (partially) observed by three calibrated perspective cameras. However, the $(5_{3},5_{3},3_{3})$ configuration and the 5pt+P3P solver were only discussed from a theoretical point of view, aiming to report the number of solutions to this problem and not its practical performance. The 5pt+P3P solver is used by some researchers [38] and practitioners in their applications [49]. Nister et al. [39] showed that the 5pt+P3P solver performs better than their dedicated 4p3v solver. Still, the 5pt+P3P solver is not even used as a baseline for a comparison in papers that study the relative pose problem for three calibrated cameras [15, 21]. To our knowledge, the performance of this solver on real data and within state-of-the-art RANSAC frameworks in the context of the 4p3v problem has not been extensively studied. Our paper fills this gap in the literature.
@@
Motivated by the efficient 5pt+P3P solver and the minimal $(5_{3},5_{3},3_{3})$ configuration, which, compared to the $(4_{4},4_{4},4_{4})$ configuration, leads to significantly simpler systems of polynomial equations, we next describe two groups of novel solvers to the calibrated 4p3v problem. These solvers efficiently solve the $(4_{4},4_{4},4_{4})$ configuration by generating a virtual point correspondence in the first two views, and thus solving it using the 5pt+P3P solver.
@@
4p3v(M) solver:  Our first solver is based on a simple observation: If we fix the $5^{\text{th}}$ point in the first view to be the mean $\mathbf{m}^{1}$ of three points $\left\{\mathbf{x}_{i}^{1},\mathbf{x}_{j}^{1},\mathbf{x}_{k}^{1}\right\},\;i,j,k\in\left\{1,\dots,4\right\}$ in this view, then the mean point $\mathbf{m}^{2}$ of the corresponding three points $\left\{\mathbf{x}_{i}^{2},\mathbf{x}_{j}^{2},\mathbf{x}_{k}^{2}\right\}$ in the second view usually has a small epipolar error w.r.t. the ground truth relative pose. Thus $\mathbf{m}^{1}\leftrightarrow\mathbf{m}^{2}$ is usually a good approximation of a correct correspondence.
@@
This can be considered as a surprising fact, since 4 (or actually 3) points in two views define an infinite number of camera poses that can observe these points. However, the reason this mean-point strategy works in practice comes from three simple facts and observations. (1) To generate a good correspondence, we only require that the point in the second view be reasonably close to the epipolar line defined by the mean point $\mathbf{m}^{1}$ in the first view, i.e., the 2D point does not need to correspond to one particular 3D point with a given depth. (2) The epipolar line defined by $\mathbf{m}^{1}$ passes through the triangle defined by the corresponding three points $\left\{\mathbf{x}_{i}^{2},\mathbf{x}_{j}^{2},\mathbf{x}_{k}^{2}\right\}$ in the second image. Therefore, the maximum distance of $\mathbf{m}^{2}$ in the second image from the epipolar line is bounded by the maximum distance of $\mathbf{m}^{2}$ from $\left\{\mathbf{x}_{i}^{2},\mathbf{x}_{j}^{2},\mathbf{x}_{k}^{2}\right\}$. (3) Three 3D points define a plane. If this plane is not observed under very oblique and different angles in two views, then the point on the 3D plane that projects on the mean point in the first image usually projects close to the mean point $\mathbf{m}^{2}$ in the second image. We support this observation by experiments on a large amount of synthetic and real data (cf. Sec. 4, mean-point strategy).

### v3

@@ Abstract
We study challenging problems of estimating the relative pose of three cameras and propose novel efficient solutions to the configurations (1) of four points in three calibrated cameras (the 4p3v problem), and (2) of four points in three cameras with unknown shared focal length (the 4p3vf problem). Our solutions are based on the simple idea of generating one or two additional virtual point correspondences in two views by using the information from the locations of the input correspondences. We generate such correspondences using a very simple and efficient strategy, where the new points are the mean points of three corresponding input points. The new solvers are efficient and easy to implement, since they are based on existing efficient minimal solvers, i.e., the well-known 5-point and 6-point relative pose solvers and the P3P solver. Extensive experiments on real data show that our solvers achieve state-of-the-art results. We also present a simple network that can improve the precision of the mean-point correspondences, showing the potential to learn better virtual point correspondences.
@@ 1 Introduction
Camera geometry estimation is crucial in many computer vision applications, e.g., visual navigation [48], Structure-from-Motion [52], augmented reality [5], self-driving cars [17], and visual localization [47]. Due to noise and outliers in the input correspondences, the predominant way for camera geometry estimation is to use a hypothesis-and-test framework, e.g., RANSAC [15, 8, 45, 2]. For RANSAC-like methods, using as few (ideally the minimal number of) correspondences as possible for estimation is important since the number of RANSAC iterations (and, thus, its run-time) grows exponentially with the number of correspondences required for the model estimation.
@@
Figure 1: Visualization of the 4p3v problem and our solution that is based on generating new virtual correspondence(s) between two views. This is done using coordinates of input point correspondences. Then the 4p3v problem is solved by existing efficient minimal solvers, i.e., the 5pt [34] and the P3P solvers [39].
@@
Minimal camera geometry problems often result in complex systems of polynomial equations. Efficient algebraic methods helped to solve many previously unsolved problems  [55, 4, 28, 25, 24, 54]. Still, they fail to generate efficient and/or numerically stable solutions for some problems. In this paper, we study such challenging problems of estimating the relative pose of three cameras. These problems have received attention for a long time [18, 44, 30, 33, 1]. However, due to their complexity, they are still not considered fully solved. There are no efficient and practical solutions for most of the minimal configurations of point and/or line correspondences [22]. One such configuration that is particularly interesting is the notoriously difficult configuration of four points in three views [44, 34]. This configuration is minimal for cameras with an unknown shared focal length, i.e., the 4p3vf problem, and provides one more constraint than minimal for calibrated cameras, i.e., the 4p3v problem.
@@
State-of-the-art algebraic and numerical methods are known to fail in generating efficient and numerically stable solutions to the 4p3v and 4p3vf problems. The existing methods for solving these problems are either too slow for practical applications [7, 9] or are only approximate [19, 35]. By solving only for one (a few) solutions from the 272 solutions of the 4p3v problem [19], and by discretely sampling the space of potential solutions [35], the existing 4p3v methods can often fail, i.e., the returned solution can be, in general, arbitrarily far from the geometrically correct solution. To decrease the failure rate, both methods [19, 35] require a lot of tuning and are not easy to re-implement. [^1] For the solver proposed in [35], there is no publicly available implementation. The publicly available implementation of the solver from [19] is quite complex and requires a non-negligible effort to run.
@@
In this paper, we propose a novel approach for solving the 4p3v and 4p3vf problems. Our solutions are based on the simple idea of generating new approximate point correspondence(s) between two of the three views. [^2] Note that similarly to the state-of-the-art solutions [35, 19], our solutions are only approximate. However, as we show in the experiments, they provide good initialization for local optimization [8] and outperform [19]. Such approximate correspondences are generated using only the locations of the original input point correspondences, without any information about the image itself (e.g., appearance or local features). Consequently, the new correspondences do not need to correspond to any physical 3D points in the scene. Thus, we call them virtual correspondences. Using virtual correspondences, we can efficiently solve the 4p3v and 4p3vf problems by first estimating the relative pose of two cameras from five/six correspondences using efficient 5pt/6pt solvers [34, 56], and then registering the third camera using a P3P solver [39]. We call these combinations the 5pt+P3P and 6pt+P3P solvers.
@@
Based on this idea, we propose a group of novel solvers for the 4p3v and 4p3vf problems. These solvers, called M-based solvers (4p3v(M),4p3vf(M)), use the mean points of three corresponding points detected in two views and, potentially, points in their vicinity ((M±$\pm\delta$)-solvers) as new virtual point correspondences. To compensate for noise in these virtual correspondences, our solvers refine the solutions on the original four points in three views using just a few iterations of Levenberg-Marquardt refinement. While conceptually very simple and efficient, the novel solvers achieve state-of-the-art results on real data.
@@
  • For the well-known challenging 4p3v problem for calibrated cameras, we propose novel M-based solvers. These solvers generate an additional virtual point correspondence(s) in two views as the mean points of three corresponding points and refine the approximate solution on the original four points in three views. The new solvers achieve state-of-the-art results in terms of accuracy on real data. Compared to state-of-the-art 4p3v solvers [19, 35], which are non-trivial and difficult to re-implement such that they are numerically stable and fast, our new solutions can be easily implemented using existing efficient implementations of the 5pt solver [34] and the P3P solver [39]. The source code of our solvers will be publicly available.
  • We provide efficient solutions to the 4p3vf problem for cameras with an unknown shared focal length. Our novel solvers generate for each instance two virtual correspondences to solve the problem via the efficient 6pt [34] and the P3P [39] solvers. Our solutions are significantly faster than the existing homotopy-continuation solutions [7, 9]. Our solvers show the potential of virtual correspondences to be applied to other camera geometry problems.
  • We present preliminary results for a simple network that can improve the precision of the mean-point correspondences. While our current versions of the learning-based (L-based) solvers do not provide sufficient improvement of virtual correspondences to produce a visible improvement after the refinement on all four correspondences, the proposed network shows the potential to learn better virtual point correspondences.
  • To the best of our knowledge, we are the first to extensively evaluate solutions to the 4p3v and 4p3vf problems on a large variety of real-world scenes and within state-of-the-art RANSAC frameworks, and to compare them to the baseline minimal 5pt+P3P and 6pt+P3P solvers on such data.
@@ 2 Related work
For three uncalibrated cameras, 6 point correspondences are necessary to estimate the trifocal tensor, with a solution known for a long time [43, 57]. Solutions to three minimal combinations of points and lines are presented in [36]. The minimal configuration using 9 lines is much more challenging and was solved only recently  [27]. However, the final solver is far from practical since it runs for 17.8s.
@@
For calibrated cameras, the configuration that attracts most of the attention is the configuration of four points in three views (the 4p3v problem). Note that this is not a minimal configuration since it generates 12 constraints for 11 degrees-of-freedom (DoF) (see also Section 3). The 4p3v problem is known to be extremely difficult to solve. Several papers present mostly theoretical results [30, 33, 1]. For four triplets of exact points without noise, it is shown that the 4p3v problem has, in general, a unique solution [18, 44].
@@
To the best of our knowledge, there are only two reasonably efficient solutions to the 4p3v problem reported in the literature. The first solver [35] is based on a one-dimensional exhaustive search. It performs a sweep of a $10^{th}$-degree curve of possible epipoles. For each potential epipole, it computes the relative pose of two cameras, registers the third camera using three triangulated points, and finally extracts the solution minimizing the reprojection error of the fourth point in the third view. Evaluation of the solver on one potential epipole is fast. Yet, to obtain reasonable precise and stable results, usually, 1,000 candidates need to be evaluated and even then, refinement at multiple local minima is required to improve the precision. The runtimes reported for this solver were $1-12ms$ depending on the number of points searched. There is no publicly available implementation for this solver and it is not easy to re-implement. As such, the literature does not compare against the solver in experiments. As an upper bound of the performance of [35], we compare with an oracle version using the true epipole in the supplementary material (SM).
@@
The second efficient solver to the 4p3v problem was published only recently [19]. In this paper, the authors first transform the 4p3v problem into a minimal problem by considering a line passing through the last correspondence in the third view. The resulting system of equations is solved using an efficient Homotopy continuation (HC) method [14, 53]. To avoid computing large numbers of spurious solutions, an MLP-based classifier is trained. For a given problem $p$, it selects one or several starting problem-solution pairs (so-called anchors), such that the geometrically meaningful/correct solution of $p$ can be obtained by HC starting from this anchor. This strategy is fast, running $16.3\mu s$ on average per solution. However, it has a high failure rate. The success rate of the 4p3v solver reported in [19] on two test datasets and data without noise is $26.3\%$. [19] do not show results for a real scenario, i.e., a RANSAC-like framework with noisy data. Providing such an evaluation, we show that our much simpler solvers outperform [19].
@@
Solutions to the 4p3v problem for orthographic and scaled orthographic views were presented in [59, 32]. In [32], the author suggested an iterative approach for updating to perspective views, but reported results only on a few synthetic instances. According to our own experiments, the update does not work on real data with general perspective cameras, even after spending months on this issue.
@@
Minimal configurations of points and lines in three calibrated views were studied in [13, 22, 12], aiming to classify and derive the number of solutions for different configurations. Solutions to two minimal configurations of combinations of points and lines (Chicago and Cleveland), were proposed in [14] and solved using a HC method [53]. Due to their complexity, the solvers are not practical,
Recently, the HC method was used to solve the 4p3vf problem for cameras with an unknown shared focal length [7, 9]. The running times of the CPU variants of the proposed solvers range from $250ms$ to $1456ms$. Efficient GPU implementations run $16.7ms$ to $154ms$. These times are still too slow for practical applications. Due to slow run-times and their dependency on a GPU, we are not comparing with the GPU solvers on real data. A GPU HC method was also used to solve minimal problems of four points/six lines in three views for a generalized camera in [11].
@@
In our solutions, we generate virtual correspondences. Virtual matches are also used in the literature on affine correspondences (AC) [38, 40, 41, 3]. There, points are sampled based on the affine feature geometry to generate point correspondences from affine ones. In our scenario, we are only given point correspondences, without associated feature geometry, and we predict additional point matches.
@@ 3 Estimating the relative pose of three cameras
Let $N$ cameras observe a set of 3D points $\mathcal{P}$. For each point $\mathbf{P}\in\mathcal{P}$, let $C_{\mathbf{P}}>1$ be the number of cameras that observe it. A necessary condition for a relative pose problem of $N$ calibrated cameras to be minimal is [14]
@@
Let $S_{m}$ denote a sample of $m$ 3D points and let $S_{m}^{k}$ denote a subset of $S_{m}$ with cardinality $k$. A configuration of points in $N=3$ views that satisfies the constraint (1) of a minimal problem is three points visible in all three cameras and two additional points visible in two of the three cameras. We will call this configuration ${(S_{5},S_{5},S_{5}^{3})}$.
The configuration of four points visible in all three cameras, i.e., the configuration ${(S_{4},S_{4},S_{4})}$, generates an over-constrained problem. In this case, we have one more constraint than DoF, i.e., in Eq. (1), we have $12>11$. A minimal solution would need to drop one constraint, e.g., by considering only a line passing through one of the points in the third view [19] or by considering a “half” point correspondence. Since, in practice, we always have full correspondences and sampling one less point in one view leads to an under-constrained problem, the configuration ${(S_{4},S_{4},S_{4})}$, is usually considered “minimal”.
For cameras with an unknown common focal length, we have one more DoF. This means that, for $N=3$, the right-hand side of equation (1) becomes $18-6=12$, resulting in ${(S_{6},S_{6},S_{6}^{3})}$ and ${(S_{4},S_{4},S_{4})}$ being minimal configurations.
@@ 3.1 Calibrated cameras
In this section, we describe solutions for three calibrated cameras. We start with one baseline minimal solution for the minimal ${(S_{5},S_{5},S_{5}^{3})}$ configuration, followed by novel solutions for the “minimal” ${(S_{4},S_{4},S_{4})}$ configuration.
@@
5pt+P3P solver:  The 5pt+P3P solver first estimates the relative pose of two cameras from 5 image point correspondences using the efficient 5pt solver [34]. Next, the three points visible in all three views are triangulated. Finally, the third camera is registered using the three 2D-3D point correspondences and the well-known efficient P3P solver [39].
@@
This solver is straightforward and it is based on efficient existing solvers [34, 39]. This solver appears in the literature [13, 34, 46, 35]. Nister et al. [35] showed that the 5pt+P3P solver performs better than their dedicated 4p3v solver. However, in the most recent works [14, 19] that study the relative pose problem for three calibrated cameras, the 5pt+P3P solver is not discussed and is not used as a baseline for comparison. To our knowledge, the performance of this solver on real data and within state-of-the-art RANSAC frameworks in the context of the 4p3v problem has not been extensively studied. Our paper fills this gap in the literature.
@@
Motivated by the efficient 5pt+P3P solver and the minimal ${(S_{5},S_{5},S_{5}^{3})}$ configuration, which, compared to the ${(S_{4},S_{4},S_{4})}$ configuration, leads to significantly simpler systems of polynomial equations, we next describe novel solvers to the calibrated 4p3v problem. They efficiently solve the ${(S_{4},S_{4},S_{4})}$ configuration by generating a virtual point correspondence in the first two views, and solving the resulting ${(S_{5},S_{5},S_{5}^{3})}$ problem using the 5pt+P3P solver.
@@
4p3v(M) solver:  Our first solver is based on a simple observation: If we fix the $5^{\text{th}}$ point in the first view to be the mean $\mathbf{m}^{1}$ of three points $\left\{\mathbf{x}_{i}^{1},\mathbf{x}_{j}^{1},\mathbf{x}_{k}^{1}\right\},\;i,j,k\in\left\{1,\dots,4\right\}$, in this view, then the mean point $\mathbf{m}^{2}$ of the corresponding three points $\left\{\mathbf{x}_{i}^{2},\mathbf{x}_{j}^{2},\mathbf{x}_{k}^{2}\right\}$ in the second view usually has a small epipolar error w.r.t. the ground truth relative pose. Thus $\mathbf{m}^{1}\leftrightarrow\mathbf{m}^{2}$ is usually a good approximation of a correct correspondence.
@@
This can be considered a surprising observation, since 4 (or actually 3) points in two views define an infinite number of camera poses that can observe these points. However, the reason this mean-point strategy works in practice comes from several simple facts and observations. (1) To generate a good correspondence, we only require that the point in the second view be reasonably close to the epipolar line defined by the mean point $\mathbf{m}^{1}$ in the first view, i.e., the 2D point does not need to correspond to one particular 3D point with a given depth. [^3] By fixing a point in one view, we are defining an epipolar line in the second view. Any of the points on this line (corresponding to 3D points with different depths) is in correspondence with the point in the first view. (2) The epipolar line defined by $\mathbf{m}^{1}$ passes through the triangle defined by the corresponding three points $\left\{\mathbf{x}_{i}^{2},\mathbf{x}_{j}^{2},\mathbf{x}_{k}^{2}\right\}$ in the second image. Thus, the maximum distance of $\mathbf{m}^{2}$ in the second image from the epipolar line is bounded by the maximum distance of $\mathbf{m}^{2}$ from $\left\{\mathbf{x}_{i}^{2},\mathbf{x}_{j}^{2},\mathbf{x}_{k}^{2}\right\}$ (for a proof, see SM). (3) For practical applications, when used in RANSAC, it is not necessary that each triplet $\left\{\mathbf{x}_{i}^{2},\mathbf{x}_{j}^{2},\mathbf{x}_{k}^{2}\right\}$ generates a good correspondence $\mathbf{m}^{1}\leftrightarrow\mathbf{m}^{2}$. Samples with a high level of noise in the mean-point correspondence are filtered inside RANSAC. [^4] This property was also used in the HC solver [19], which completely fails for many samples. These samples are filtered within RANSAC. On a large number of different scenes, we observed that even if some image pairs have triplets of points that generate very noisy mean-point correspondences, there are usually enough triplets for which the noise in $\mathbf{m}^{2}$ is reasonably small to provide good estimates. (4) Four point correspondences in two views usually fix the space of possible poses such that the $5^{th}$ correspondence, even if noisy, often generates a pose that is not very far from the ground truth pose. Such a pose is usually sufficient as an initialization of non-linear optimization on the original four points in three views and subsequent local optimization on detected inliers. We support our observations by experiments on a large amount of synthetic and real data (see Sec. 4 and SM).